# LLMs and the Human Condition

### v2

@@ Abstract
Theory based AI research has had a hard time recently and the aim here is to propose a model of what LLMs are actually doing when they impress us with their language skills. The model integrates three established theories of human decision-making from philosophy, sociology, and computer science. The paper starts with the collective understanding of reasoning from the early days of AI research - primarily because that model is how we humans think we think, and is the most accessible. It then describes what is commonly thought of as "reactive systems" which the position taken by many philosophers and those contemporary AI researchers who think about it. The third component to the proposed model is from sociology which, although not flattering to our modern ego, it does provide an explanation to a puzzle that has for many years occupied those of us working on conversational user interfaces. Today of course questions of what it means to be intelligent and how to do it have been superseded by a belief that Neural Net architectures provide the answers. If there is discussion of what it is to be intelligent or the nature of representation and language, it is based on discredited theory. There is an awful lot of marketing and "I want to believe" at the moment, and although the results the machine learning people are getting are certainly impressive, it is clear they are not experts on the problem their machines solve. This paper provides an explanation for the success of Large Language Models that fits with what we know about the nature of language and the way humans make decisions.
@@
In AI at one point researchers were classified as “neats” or “scruffies”. The neats worked from theory, while the scruffy researchers took an extreme programming approach, making things that did something and then tinkering to improve it. The theory people had trouble making things work, but learnt lots about their theories. Often, as is the case with Brooks seminal paper “Intelligence without Representation” Brooks (1991), the scruffies shunned theory, claiming to be just doing engineering. This is a neats paper but perhaps, if we understand what is happening when a transformer assembles a LLM from all the text in the world, we can suggest a better way forward for those ’just’ engineering.
In what follows I first recap good old fashioned AI (GOFAI) and then describe what Brooks did, and what philosophers (and psychologists) have said. I then introduce a third mechanism that was touched on in the 1970s in AI research but is an idea that is hundreds of years old in what is now sociology. Finally I show how the model can be applied to a major issue in Natural Language Understanding, namely the phenomena of apparent “mind reading”.
@@ 1 The problem: people read minds
Working on NLU, one sooner or later comes across the phenomena in which humans appear to do “mind reading” Bar-On (2021). In what follows the focus is one example but the phenomena itself is extremely common. The solution is often assumed to be a better (and more detailed) representation of the context and from that perspective, the proposed solution is to represent context, not as things, but as practices.
@@
1 Child:
I’m hungry
2 Mother:
Did you do a good job of your geography homework?
3 Child:
Yeah.
4 Child:
What’s to eat?
5 Mother:
Let me read it.
6 Mother:
What is the capital of Brazil?
7 Child:
Rio de Janeiro.
8 Mother:
Think about it.
9 Child:
It’s Brasilia.
10 Child:
Can I eat now?
11 Mother:
I’ll let you have something later.
12 Mother:
What is the capital of Venezuela?
13 Child:
Caracas.
14 Mother:
Fine.
15 Child:
So what can I eat?
16 Mother:
You want some cereal?
17 Child:
Sure.
All English speakers by definition Tomasello (2008) will follow what is happening in this conversation and feel there is nothing extraordinary about it. The problem for a reference model of language understanding is that the mother says nothing related to the child’s utterances until line 11. There is no overlap of semantic content - none at all. Of course as English speakers we can all say what both the mother and the child want.
This has historically lead researchers interested in conversational agents off into the realm of recognition of intent and modelling theory-of-mind (ToM) Allen et al. (1995). Classically this would be modelled in terms of the beliefs, desire and intentions of an agent, the theory being that a rational agent will do what it believes is in its interests Cohen and Levesque (1990), all modelled in terms of good old fashioned symbolic representations. Rather than falling back on GOFAI notions of what intelligence is, perhaps there is another way to read minds.
@@ 2 Machines
Good old fashioned AI was based on a reference model of meaning in which it is assumed symbols such as the word ’cat’ has meaning by referring to a thing in the world, namely the set of all cats. This is generally how people see language as working, and the early AI researchers saw the problem as one of tidying the hap-hazard oddities found in natural languages. For centuries there has been the idea that meaning bottoms out at some kind of semantic primitives Ishiguro (1990) and in this tradition Roger Schank proposed that there was a workable set of primitive verbs or ‘acts’ from which the meaning of a sentence could be based Schank (1972). Examples are PTRANS for “physical transfer” and MTRANS for mental or knowledge transfer, and that machines could translate natural descriptions into (representations of) meaning and figure out consequences from there. The universal primitives approach to AI didn’t really work but there was also the idea that a formal system of symbol manipulations might map isomorphically onto events in the world. In “Godel, Escher, Bach”, Hofstader Hofstader (1979) gives a beautiful example in the form of his pq- system which introduces the idea of meaning in formal (symbol) systems via isomorphism. Although the symbols themselves are meaningless, they happen to map onto the number of apples in a box as you add more, or take them out. The meaning of “-p--q---” comes from its location in a “structure” that is isomorphic with the apples in the box after putting in another two – no primitives involved. Between computers being good at this kind of symbol manipulation, and Schank’s observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things. The question is which columns to have, and how things in the world get allocated to them.
@@
There turns out to be a wide range of tasks that fail when handled this way with a computer. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many worms and rotten ones there are. Often the issue is speed, so if the idea is to map a room before you vacuum it, things that move like cats and people tend to re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, identified by flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world.
The feeling in the late 1980s was that choosing a system of symbols - a set of signs (e.g. “CAT”) and signified (e.g. the set of all domestic cats, and wild cats that someone might domesticate, and …) - was problematic. What is more, there was a better way. Rather than using sense-data to form a representation of the world, then reasoning about it before acting, functional systems could be built by connecting sensing to acting. Roomba vacuum cleaners do a random walk of a room to clean it and avoid chairs (and cats) by touching them and turning away. Since 1990 there has been a glacial shift in the computer science collective understanding of the problem and today there is wide spread acknowledgement that the nature of meaning is inseparably linked to our embodiment in the world Wooldridge (2023).
@@
How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence. Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick. From a philosopher’s perspective, the system is more interesting when considered as just a part of the agent-environment system. ELIZA worked in a particular environment in exactly the same way as Roombas work in a particular environment and, the claim is, in exactly the same way we humans make most of our decisions about what to do next.
@@ 3 Humans
Deciding if I should buy that bottle of whiskey might involve reasoning with symbols representing money, but getting something to eat when I know where the refrigerator is - that is perhaps literally a no brainer, not only available to the cat, but also to me. The radical enactivists Ward et al. (2017) point out how agents and their environment are often set up so that the environment triggers; the environment “brings on”, or the agent “directly perceives” something that causes the agent to behave in an appropriate way with no need for intermediate representation or indeed thinking. Like the cat, when hungry I automatically move to the refrigerator where the handle “calls out” to me to open the door. Having opened the door, a packet of gnocchi catches my eye (because I am hungry) and picking up the packet, the pan-handle and tap guide my next actions. Like the insect crossing a pebble beach [^1] I recall this example from undergraduate lectures but can’t find a citation. the behaviour looks complex. Indeed we might be tempted for whatever reason to ascribe beliefs and desires to the insect, but what unfolds is a “trajectory” produced by the agent but directed by the environment. Looking back on my actions I will of course explain them in terms of beliefs and goals, but in the act of doing them, the mechanism does seem to be reactive and hence completely uninteresting to computer scientists.
@@
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment. It seems safe to assume other creatures do not. Although beavers might plan to build a dam, termites no doubt use their “insect level” intelligence to build termite mounds. The question becomes how much of the human environment has been constructed with our “termite intelligence”.
@@
Recently we have been talking with sociologists who look at human decision making at the macro level. The strong reductionist position is alive and well in sociology taking the line that collective action is simply made up from the behaviour of individuals making their own decisions - basically that sociology is just psychology on a massive scale in the same way as drunk physicists will claim chemistry is just physics. Social simulation based on the behaviour of individuals is of course entirely possible with today’s computers. There are however problems with this in that collective action does not fit with rational self interest of individuals. Why did Britain leave the EU? And why does Germany want to go to war with a nation that 2 years ago was a key trading partner? At the other end of the “structure and agency” debate wikipedia in sociology is the idea that society consists essentially of structures such as institutions, traditions, and norms, and that individuals are merely acting out their role in those structures. Post The Enlightenment we collectively like to think we can reason about things and make rational decisions that are better than the status quo and it is of course not very flattering to be told you are merely a cog in a machine. A key finding of AI research over the last 70 years is of course that rationality is not all it is cracked up to be. We humans do situated action using “insect level” intelligence in a benign environment, but setting up that benign environment is well beyond the capabilities of me or any other individual. A group of the esteemed homo economicus, sitting on a desert island are not going to invent agriculture, houses, trade, utility companies, the stock market and war. These are handed down to us from those who came before. We are also raised to know how to use them.
The existence of these institutions looks designed of course but, the theory goes, they are actually evolved structures of successful societies. Rather than smart people having a meeting and inventing the stock-market, the stock-market, like the human eye, evolved from what came before. The stock-market might have been set up by a committee, but concepts such as trade and money already existed, as did the process of collective decision-making by committee. Whatever the mechanism of change, the current institutions have been selected for by evolutionary pressures. Society is an “ecology of practices” where the practices are the thing that survive and people are merely the actuators for roles. As our techie once said, “Cogito, ergo, I’m a cog.” Leon (1986).
@@
The society’s strategies may or may not accord with the survival strategies of individual humans of course - war being a classic example - and the survival strategies of society may not be particularly accessible to an individual’s self reflection. Perhaps, rather than termite mounds being indicators of distributed intelligence in termites, cathedrals are indicators of “termite intelligence” in humans?
The proposal is that we do have the ability to do rational choice, but it is constrained between the realities of being embodied in a physical world, and those of being born into an operational society. The strong linguistic relativism argument is that l’sign is indeed arbitare Culler (1976) but the things all humans want to talk about are within these contingencies, contingencies that are shared across all cultures. The suggestion is that animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures. The unique human skill is the practice of “thingifying” the stuff around us, giving it a label, and then reasoning symbolically with the label. Language is both a window on how that is done, and a mechanism for passing it on.
@@ 4 A ’pragmatics first’ model of NLU
Whereas sociologists are interested at decision making at the macro level, the interest here is the mechanism for decision making at the agent level. Our aim is to find a plausible model of what goes on in heads. The position taken here is that language is indeed a window on the mind, but only the conscious part. We catch balls, and we catch a bus. We do not talk about the mechanism of eye-hand coordination directly, and we don’t “see” the intricate workings of the queue at the bus stop. For those of us interested in machines using language, we need to stop thinking about the meaning of words, and instead focus on the practices in which language is used to do things.
@@
Of course in the example this does not happen. The child’s environment (i.e. Mother) does something different to what is expected at step two. In the child’s repertoire of practices however, mother’s utterance is a fit with the practice of checking homework. The child then has a decision to make: continue with the practice of getting something to eat, or to switch the PiP to checking homework. It is here that the child gets to do “purposive human action”, not about his choice of words but about his choice of practice to perform. This level of decision is not linguistic - it is not about syntax and semantics, but about all sorts of non-linguistic but very pragmatic factors such as the power relation between mother and child.

### v3

@@ Abstract
Theory based AI research has had a hard time recently and the aim here is to propose a model of what LLMs are actually doing when they impress us with their language skills. The model integrates three established theories of human decision-making from philosophy, sociology, and computer science. The paper starts with the collective understanding of reasoning from the early days of AI research - primarily because that model is how we humans think we think, and is the most accessible. It then describes what is commonly thought of as "reactive systems" which is the position taken by many philosophers and indeed many contemporary AI researchers. The third component to the proposed model is from sociology and, although not flattering to our modern ego, provides an explanation to a puzzle that for many years has occupied those of us working on conversational user interfaces.
@@
In AI at one point researchers were classified as “neats” or “scruffies”. The neats worked from theory, while the scruffy researchers took an extreme programming approach, making things that did something and then tinkering to improve it. Theory was generally based on a reference model of semantics in which symbols such as the word ’cat’ has meaning by referring to a thing in the world. Deciding if that phenomena on the end of the bed is a member of the set of all cats might be problematic but, having done that, the expression “the cat” simply refers to something in the world, and it is a cat because that is in the set of things called cats. The symbol grounding problem Harnad (1990) is a problem, but it turns out to be the wrong problem. In contrast the scruffies generally didn’t bother with theory and indeed Brooks in his seminal paper “Intelligence without Representation” Brooks (1991) claimed to be “just doing engineering” and disavowed any influence from Heidegger. The aim here is to look again at the theory, but perhaps if we understand what is happening when a transformer assembles a LLM from all the text in the world, we can suggest a better way forward for those ’just’ engineering.
The paper starts with a recap of good old fashioned AI (GOFAI) and then describes what Brooks and many philosophers (and psychologists) have said. It then introduces a third mechanism that was touched on in the 1970s in AI research but which is an idea that is hundreds of years old in what is now sociology. Finally the model being developed is applied to a major issue in Natural Language Understanding (NLU) namely the phenomenon of apparent “mind reading”.
@@ 1 The problem: people read minds
Working on NLU, one sooner or later comes across a phenomenon in which humans appear to do “mind reading”. In what follows the focus is one example but the phenomenon itself is extremely common. The solution is often assumed to be a better (and more detailed) representation of the context and from that perspective, the proposed solution is to represent context, not as things, but as practices.
@@
1 Child: I’m hungry
2 Mother: Did you do a good job of your geography homework?
3 Child: Yeah.
4 Child: What’s to eat?
5 Mother: Let me read it.
6 Mother: What is the capital of Brazil?
7 Child: Rio de Janeiro.
8 Mother: Think about it.
9 Child: It’s Brasilia.
10 Child: Can I eat now?
11 Mother: I’ll let you have something later.
12 Mother: What is the capital of Venezuela?
13 Child: Caracas.
14 Mother: Fine.
15 Child: So what can I eat?
16 Mother: You want some cereal?
17 Child: Sure.
All English speakers by definition Tomasello (2008) will follow what is happening in this conversation and feel there is nothing extraordinary about it. The problem for a reference model of language understanding is that the mother says nothing related to the child’s utterances until line 11. There is no overlap of semantic content - none at all. Of course, as English speakers, we can all say what both the mother and the child want.
This has historically lead researchers interested in conversational agents off into the realm of recognition of intent and modelling theory-of-mind (ToM) Allen et al. (1995). Classically this would be modelled in terms of the beliefs, desire and intentions of an agent, the theory being that a rational agent will do what it believes is in its interests Cohen and Levesque (1990), all in terms of good old fashioned symbolic representations.
@@ 2 Machines
The reference model of meaning is generally how people see language as working, and the early AI researchers saw the problem as one of tidying the hap-hazard oddities found in natural languages. For centuries there has been the idea that meaning bottoms out at some kind of semantic primitives Ishiguro (1990) and in this tradition Roger Schank proposed that there was a workable set of primitive verbs or ‘acts’ from which the meaning of a sentence could be based Schank (1972). Examples are PTRANS for “physical transfer” and MTRANS for mental or knowledge transfer, and that machines could translate natural descriptions into (representations of) meaning and figure out consequences from there. The primitives approach to AI didn’t really work but there was also the idea that a formal system of symbol manipulations might map isomorphically onto events in the world. In “Godel, Escher, Bach”, Hofstader Hofstader (1979) gives a beautiful example with his pq- system which is a set of syntactic transforms on strings of ’p’, ’q’, and ’-’. The system produces an infinite set of strings including “-p--q---” and “--p--q----” but not “-p---q--”. The number of ’-’s in produced strings just happens to map onto the number of apples in a box as you add more, or take them out. The symbol ’p’ seems to mean “plus” and ’q’, “equals” because they appear in the system of productions in an isomorphic relationship with our system of adding and subtracting. They have no meaning in themselves without the system in which they participate. Rather than meaning being built up by combining primitive features with innate meaning, the “primitives” get meaning from the structure in which they occur.
@@
Computers are good at this kind of syntactic symbol manipulation, and between that and Schank’s observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things. It turns out there is a wide range of tasks that fail when handled this way with a computer. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many worms and rotten ones there are. Often the issue is speed, so if the idea is to map a room before you vacuum it, things that move like cats and people tend to re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, identified by flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world.
@@
The feeling in the late 1980s was that no matter how powerful computers might become in the future, simply getting the data for symbolic representations was going to be problematic. There were projects set up to encode all the data in the world, the CYC project Lenat et al. (1986) being the most explicit, but the scruffies were making progress and perhaps there was a better way. Rather than using sense-data to form a representation of the world, then reasoning about it before acting, functional systems could be built by connecting sensing to acting. Roomba vacuum cleaners do a random walk of a room to clean it and avoid chairs (and cats) by touching them and turning away. Since 1990 there has been a glacial shift in the computer science collective understanding of the problem and today there is wide spread acknowledgement Wooldridge (2023) that the nature of meaning is inseparably linked to our embodiment in the world. How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence. Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick. From a philosopher’s perspective, the system is more interesting when considered as just a part of the agent-environment system. ELIZA worked in a particular environment in exactly the same way as Roombas work in a particular environment and, the claim is, in exactly the same way we humans make most of our decisions about what to do next.
@@ 3 Humans
Deciding if I should buy that bottle of whiskey might involve reasoning with symbols representing money, but getting something to eat when I know where the refrigerator is - that is perhaps literally a no brainer. The radical enactivists Ward et al. (2017) point out how agents and their environment are often set up so that the environment triggers; the environment “brings on”, or the agent “directly perceives” something that causes the agent to behave in an appropriate way with no need for intermediate representation or indeed thinking. Like the cat, when hungry I automatically move to the refrigerator where the handle “calls out” to me to open the door. Having opened the door, a packet of gnocchi catches my eye (because I am hungry) and picking up the packet, the pan-handle and tap guide my next actions. Like the insect crossing a pebble beach [^1] I recall this example from undergraduate lectures but can’t find a citation. the behaviour looks complex. Indeed we might be tempted for whatever reason to ascribe beliefs and desires to the insect, but what unfolds is a “trajectory” produced by the agent but directed by the environment. Looking back on my actions when cooking gnocchi I will of course explain them in terms of beliefs and goals, but in the act of doing them, the mechanism does seem to be reactive. Unfortunately is is completely uninteresting to real computer scientists.
@@
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment. However it seems safe to assume other creatures do not.
@@
Recently we have been talking with sociologists who look at human decision making at the macro level. The strong reductionist position is alive and well in sociology and many take the line that collective action is simply made up from the behaviour of individuals making their own decisions - basically that sociology is just psychology on a massive scale in the same way as drunk physicists will claim chemistry is just physics. Social simulation based on the behaviour of individuals is of course entirely possible with today’s computers. There are however problems with this in that collective action does not fit with rational self interest of individuals. Why did Britain leave the EU? And why does Germany want to go to war with a nation that 2 years ago was a key trading partner? At the other end of the “structure and agency” debate wikipedia in sociology is the idea that society consists essentially of structures such as institutions, traditions, and norms, and that individuals are merely acting out their role in those structures. Post The Enlightenment we collectively like to think we can reason about things and make rational decisions that are better than the status quo and it is of course not very flattering to be told you are merely a cog in a machine. A key finding of AI research over the last 70 years is of course that rationality is not all it is cracked up to be. We humans do situated action using “insect level” intelligence in a benign environment, but setting up that benign environment is well beyond the capabilities of me or any other individual. A group of the esteemed homo economicus, sitting on a desert island are not going to invent agriculture, houses, trade, utility companies, the stock market and war. These are handed down to us from those who came before. The existence of these institutions looks designed but, the theory goes, they are actually evolved structures of successful societies. Rather than smart people having a meeting and inventing the stock-market, the stock-market, like the human eye, evolved from what came before. The stock-market might have been set up by a committee, but concepts such as trade and money already existed, as did the process of collective decision-making by committee. Whatever the mechanism of change, the current institutions have been selected for by evolutionary pressures. Society is an “ecology of practices” where the practices are the thing that survive and people are merely the actuators for roles. As our techie once said, “Cogito, ergo, I’m a cog.” Leon (1986).
@@
The society’s practices may or may not be visible to us mere mortals, and of course society’s practices may not accord with the survival strategies of individual humans – war being a classic example. Social insects like ants and bees do not lay down their lives out of love for their fellow insect, but because they are wired that way. Indeed an individual ant does not (let’s assume) have a strategy, or indeed even a practice. Instead the ant society has practices, including attacking intruders. The mechanism for implementing it is to produce ants that perform the role of soldier ants. As humans we may be tempted to claim that the ant nest has a strategy for protecting itself but that would suggest that the nest in some way has goals and representations of an enemy, war, and casualties. The notion of a practice is that it happens, with or without representations of the world or beliefs, with or without goals or desires, and with or without any sense of formulating an intention.
Figure 1: Sagrada Familia, Barcelona, vs Termite mounds in Western Australia
Some have drawn parallels between termite mounds and cathedrals and suggested that a termite colony must have some form of “distributed intelligence” that enables (rational) decision making. Instead it seems we should look at Sagrada Familia and think of it as representing “insect intelligence” in humans. Gaudi undoubtedly thought about Sagrada Familia, but he did it in terms of the established practices in which individual stone masons and accountants, priests, engineers and parishioners, all fill roles.
We do have the ability to do rational choice, and we do it with symbolic representations of things in the world. But following the strong version of the linguistic relativism argument, l’sign is arbitare Culler (1976) and what we reason about is culturally constructed. But the things all humans want to talk about are within contingencies, contingencies that are shared across all cultures. The suggestion is that we animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures. The unique human skill is the practice of “thingifying” the stuff around us, giving it a label, and then reasoning symbolically with the label. Language is both a window on how that is done, and a mechanism for passing it on.
@@ 4 A ’pragmatics first’ model of NLU
Whereas sociologists are interested at decision making at the macro level, the interest here is the mechanism for decision making at the agent level. Our aim is to find a plausible model of what goes on in heads. The position taken here is that language is indeed a window on the mind, but only the conscious part. We catch balls, and we catch a bus. We do not talk about the mechanism of eye-hand coordination directly, and we don’t “see” the intricate workings of the queue at the bus stop. For those of us interested in machines using language, we need to stop thinking about the meaning of words, and instead focus on the practices in which language is used to do things. Yes we do things with words Austin (1955), but the doing is embedded in practices.
@@
Of course in the example this does not happen. The child’s environment (i.e. Mother) does something different to what is expected at step two. In the child’s repertoire of practices however, mother’s utterance is a fit with the practice of checking homework. The child then has a decision to make: continue with the practice of getting something to eat, or to switch the PiP to checking homework. It is here that the child gets to do “purposive human action” – to reason about what to do next, The thing reasoned about is not his choice of words however, but about his choice of practice to perform. This level of decision is not linguistic - it is not about syntax and semantics, but about all sorts of non-linguistic but very pragmatic factors such as the power relation between mother and child.